OpenStudy (zuha):

Δ JKL has vertices J(10, 5), K(2, 8), and L(4, 9). It is translated right 5 units and up 4 units. What are the coordinates of J’, K’, and L’? (1 point) J' (15, 9), K' (12, 7), L' (9, 13) J' (15, 9), K' (7, 12), L' (9, 13) J' (14, 10), K' (6, 13), L' (10, 12) J' (15, 5), K' (7, 8), L' (9, 9)

OpenStudy (zuha):

please help!!

OpenStudy (loser66):

10+5 =? 2+5 =? 4+5=?

OpenStudy (loser66):

Are my questions so hard? I guide you the steps, that's the first step